OpenStudy (anonymous):

Factor completely 100x2 − 49. (6 points) (7x − 10)(7x + 10) (7x − 10)(7x − 10) (10x − 7)(10x + 7) (10x − 7)(10x − 7)

OpenStudy (anonymous):

What are the solutions of 2x2 + 3x= −8? (6 points) the quantity negative 3 plus or minus i square root of 41 all over 4 the quantity negative 3 plus or minus i square root of 55 all over 4 the quantity negative 3 plus or minus i square root of 21 all over 4 the quantity negative 3 plus or minus i square root of 15 all over 4

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the first one is C because the answer is always C also because \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (misty1212):

for the second one, use the quadratic formula can't really read those answers though, i am sure one of them is right

OpenStudy (anonymous):

thanks!!!!

OpenStudy (anonymous):

What are the solutions of 2x2 + 16x + 50 = 0? (6 points) −4 ± i −4 ± 3i −2 ± 4i −3 ± 4i

OpenStudy (misty1212):

\[2x^2+16x+50=0\] divide by 2 and start with \[x^2+8x+25=0\]

OpenStudy (misty1212):

since 8 is even, easiest to complete the square subtract 25 and get \[x^2+8x=-25\] as step one then complete the square do you know how to do that?

OpenStudy (anonymous):

oh ok i got it thank you! A right?